Potassium 40 has percentage abundance of 0.0012%. How many atoms of potassium 40 are present in 225ml of milk containing 1.65 mg K/ml?
1
Expert's answer
2020-12-07T03:18:11-0500

The total mass of potassium present in the milk sample is:

"225mL\times1.65mg/mL=371.25mg=0.37125g"


Total moles of K:


"n=\frac{0.37125g}{39.098g/mol}=9.4953\cdot10^{-3}mol" , out of which moles of K-40 are equal to


"9.4953\cdot10^{-3}mol\times\frac{0.0012\%}{100\%}=1.1394\cdot10^{-7}mol"


Number of atoms:


"1.1394\cdot10^{-7}mol\times6.022\cdot10^{23}=6.86\cdot10^{16}"


Answer: "6.86\cdot10^{16}" atoms